# On the conjugates of Christoffel words

### v4

@@ Abstract.
2010 Mathematics Subject Classification:
68R15
1. Introduction
In an article published in 1875 [C], Christoffel introduced a class of words on a binary alphabet, which now bear his name (following Berstel [B]); they were shortly after rediscovered by Smith [S]. Independently, in his 1880 work on minima of quadratic functions [M1, M2], Markoff used these words to construct certain quadratic forms, which satisfy sharp inequalities relating their minima and their discriminant. Markoff was certainly unaware of the work of Christoffel. The relation between the work of Markoff and the Christoffel words was explicitly known to Frobenius [F], who also formulated the famous conjecture on the Markoff numbers (see [A]). The theory of these words may be found in several books: [PF, L, BLRS, A, Re].
@@ 1. Introduction
1. They coincide with the elements of the free group with two generators subject to the following conditions: they are positive (that is, with no inverted letter), they are cyclically reduced, and they are part of a basis of this free group; see [OZ, KR].
@@
2. They are the “perfectly clustering words” on a two-letter alphabet; this means that the last column of the Burrows-Wheeler tableau of such a word [^1] This tableau was defined by Burrows and Wheeler in the theory of data compression, see [RR] Section 4., whose rows are the lexicographically sorted conjugates, is decreasing; see [MRS, FZ].
@@
3. They constitute the finitary version of the Sturmian (infinite) words, which are obtained by discretizing straight lines in the plane, and which are characterized by the property that for each $n$, they have exactly $n+1$ factors of length $n$. For example, a word is a conjugate of a Christoffel word if and only if all its conjugates are factors of a Sturmian word; equivalently, this word (of length $n$ say) is primitive and has exactly n1n-1 circular factors of length n2n-2; see [L], [Re, Theorem 15.3.1].
@@
4. Besides the Christoffel words, which encode the Markoff forms and their minima, their conjugates correspond to the “small values” of these quadratic forms; see [Re1].
@@
It is well known that Christoffel words are parametrized by nonnegative rational numbers. In the present article we first introduce a parametrization of the conjugates of Christoffel words, which is a finitary version of [BL]. This parametrization is based on the integer Ostrowski numeration system. It generalizes a construction which is widely used in the theory of Sturmian words, following Rauzy [R] (the “Rauzy rules”), and de Luca and Mignosi [dL] (the “standard words”). The construction is given in (7.1). Theorem 7.3 states that the whole conjugation class is constructed, and that it is independent of the chosen Ostrowski representation. As a corollary we obtain a result of Frid [Fi], which states (in some equivalent formulation) that the prefixes of a standard word are parametrized by legal Ostrowski representations; see Corollary 7.6, which appears as a noncommutative lifting of the Ostrowski numeration system.
@@
In Section 8, we study the borders (a prefix which is also a suffix) of conjugates of Christoffel words. It is well known that the length of the longest border of a word and its smallest period are simply related: their sum is the length of the word. The study of periods in words is an important matter in combinatorics on words, in particular in the theory of Sturmian words: it is known that each finite Sturmian word has a nontrivial proper period, except precisely the Christoffel words. We thus focus on conjugates of Christoffel words, and determine their longest borders. Our parametrization of the conjugates allows us to give precise statements on the form of these borders. In particular, they are themselves conjugates, or a power of them (Theorem 8.1, Corollaries 8.2 and 8.3). Note that smallest periods of conjugates of Christoffel words have been previously computed by Lapointe [La], and applied by her to the determination of normal forms, thereby allowing her to characterize conjugates within the class of Sturmian words. The set of smallest periods is also studied in [HN] and [CS].
@@
In Section 9, we give an application of our methods to notions and results due to Epifanio, Frougny, Gabriele, Mignosi and Shallit [EMSV, EFGMS]. The result of Frid, once formulated for the so-called “lazy” Ostrowski representation (a notion introduced by these authors), may be translated into a result on the paths of a certain graph, called the “compact graph”; it states that each suffix of the central word corresponding to a Christoffel word is the label of a unique path, starting from the origin, in this graph. By specializing to lengths, one obtains the “Sturmian graph”: this graph has the property that each integer from 0 to the length of the central word is the label of a unique path; see Corollaries 9.2 and 9.3, due to [EMSV, EFGMS]. As a consequence, we obtain the new result that these two graphs are naturally embedded in the tree of central words and in the Stern-Brocot tree; see Corollary 9.5, for the proof of which we use the “iterated palindromisation” of Aldo de Luca.
@@
Note that a link between lazy representations and periods of words was already established by Gabric, Rampersad and Shallit [GRS]: they determine the set of periods of each prefix of length $n$ of a characteristic Sturmian (infinite) word and they show that the cardinality of this set is equal to the sum of the digits in the lazy representation of $n$.
@@ 2. Continuant polynomials
for any $k\geq 1$ (right recursion formula). It is customary to drop the index $k$ and to write $K(n_{1},\ldots,n_{k})$ for $K_{k}(n_{1},\ldots,n_{k})$, and in particular $K()=1$. One has (for example [Co] p. 116)
(2.1) $P(n_{1})\cdots P(n_{k})=\left(\begin{array}[]{cc}K(n_{1},\ldots,n_{k})&K(n_{1},\ldots,n_{k-1})\\ K(n_{2},\ldots,n_{k})&K(n_{2},\ldots,n_{k-1})\end{array}\right),$
@@
It follows also, by transposing the product, and using the symmetry of the matrices $P(n)$, that $K(n_{1},\ldots,n_{k})=K(n_{k},\ldots,n_{1})$.
@@
(2.2) $K(n_{1},\ldots,n_{k})=K(n_{1}-1,n_{2},\ldots,n_{k})+K(n_{2},\ldots,n_{k}),$
@@ 3. Ostrowski numeration
The representation (3.2) is called legal if one has the inequalities
(3.3) $\forall i\geq 1,0\leq d_{i}\leq b_{i}.$
Among the legal representations, we distinguish two of them. We say that the representation (3.2) is greedy if it is legal and if the following condition is satisfied
(3.4) $\forall i\geq 2,d_{i}=b_{i}\Rightarrow d_{i-1}=0.$
We say that the representation (3.2) is lazy if it is legal and if, with $k=\max\{i\mid d_{i}\neq 0\}$,
(3.5) $\forall i,2\leq i\leq k,d_{i}=0\Rightarrow d_{i-1}=b_{i-1}.$
Proposition 3.1.
@@
The existence of a representation (3.2) is implicit in Ostrowski’s article [O, p.178]; (i) is stated in [D] p.83, and proved by Fraenkel, [Fr, Theorem 3] (see also [AS, Theorem 3.9.1] for a proof). Lazy Ostrowski representations were introduced by Epifanio, Frougny, Gabriele, Mignosi and Shallit in [EFGMS]; (ii) follows from their work.
@@
Lemma 3.2.
Let $\sum_{j=1}^{m}d_{j}q_{j-1}$ be a greedy Ostrowski representation. Then $\sum_{j=1}^{m}(b_{j}-d_{j})q_{j-1}\leq q_{m-1}-1$ if and only if $d_{m}=b_{m}$ and the sequence $d_{i},i=1,\ldots,m$, is alternating.
Proof.
By Proposition 3.1 (ii), $\sum_{j=1}^{m}b_{j}q_{j-1}=q_{m}+q_{m-1}-2$ since the left-hand side is a lazy representation, and is necessarily the largest one. Thus the inequality of the lemma is equivalent to $\sum_{j=1}^{m}d_{j}q_{j-1}\geq q_{m}-1$. By the proposition again, part (i) this time, this inequality is equivalent to the fact that the left-hand side is the unique greedy representation of $q_{m}-1$. But by Lemma 10.1 (i), this unique representation is the alternating one, with $d_{m}=b_{m}$. ∎
4. Conjugation
@@
Since $C^{n}$ is the identity on the words of length $n$, we have $C^{|g|}(u)=C^{r}(u)$, and this implies the last assertion. ∎
5. Morphisms
@@
(5.2) $\left(\begin{array}[]{cc}|f(g)|_{a}\\ |f(g)|_{b}\end{array}\right)=M(f)\left(\begin{array}[]{cc}|g|_{a}\\ |g|_{b}\end{array}\right).$
@@ 6. Christoffel words
The conjugation class of some Christoffel word is called a Christoffel class. It is known that in each Christoffel class, there is exactly one lower, and one upper, Christoffel word. See [L, R] for this and other properties of these words.
@@
7. Construction of the conjugates of a Christoffel word
@@
Following [BL], given a sequence of integers $d_{1},\ldots,d_{m}$ in $\mathbb{Z}$}, we define the following sequence $V_{i}=V_{i}(d_{1},\ldots,d_{m})$, of elements of $F(A)$, by
@@
The last one follows from the injectivity of the morphisms $\pi(i,j)$ (because they extend to automorphisms of the free group), hence of their product. ∎
@@
Thus by (2.1) $|V_{m}|_{a}=K(b_{1},\ldots,b_{m})$ and $|V_{m}|_{b}=K(b_{2},\ldots,b_{m})$. We have $b_{i}=a_{i}$, except for $i=1$, where $b_{1}=a_{1}-1$. Thus (i) follows, using (2.2) for the third formula, and (ii) follows at once. ∎
If the sequence $d_{1},\ldots,d_{m}$ satisfies the inequalities (3.3), then the exponents in (7.1) are all nonnegative, therefore $V_{i}(d_{1},\ldots,d_{m})\in A^{*}$. Define
(7.2) $M_{m}=V_{m}(0,\ldots,0)=\pi(b_{1},0)\circ\cdots\circ\pi(b_{m},0)(a),$
@@
The $A^{*}$-conjugation class of $M_{m}$ is equal to the set of all $V_{m}(d_{1},\ldots,d_{m})$, for all sequences $d_{1},\ldots,d_{m}$ satisfying (3.3) and precisely $V_{m}=C^{N}(M_{m})$, with $N$ as above. This class contains the two Christoffel words of Slope $S=[0,a_{1},\ldots,a_{m}]$.
@@
(7.3) $a^{-d_{1}}(\pi(b_{1},0)(a^{-d_{2}}))\cdots(\pi(b_{1},0)\circ\cdots\circ\pi(b_{m-1},0)(a^{-d_{m}})).$
@@
(7.4) $V_{m}=h^{-1}M_{m}h$
and, by (5.3) and (5.2), the algebraic length of $h$ is equal to
@@
by (2.2), since $b_{i}=a_{i}$ if $i\geq 2$, and $b_{1}=a_{1}-1$.
If the sequence $d_{1},\ldots,d_{m}$ satisfies (3.3), then $V_{m}$ is in $A^{*}$, and by Lemma 4.1, $V_{m}=C^{N}(M_{m})$, thus $V_{m}$ is in the conjugation class of $M_{m}$. Conversely, each element of this class appears, since $M_{m}$ is of length $K(a_{1},\ldots,a_{m})$ (Lemma 7.2), and since, by Proposition 3.1, each $N=0,\ldots,K(a_{1},\ldots,a_{m})-1$ has an Ostrowski representation satisfying (3.3).
We show now that the class contains a Christoffel word. Consider the sequence $d_{1},\ldots,d_{m}$ defined by $d_{m}=b_{m},d_{m-1}=0,d_{m-2}=b_{m-2}$, and so on, depending on the parity of $m$. The corresponding element $V_{m}$ is in $A^{*}$, and is equal to $\cdots\pi(0,b_{m-2})\circ\pi(b_{m-1},0)\circ\pi(0,b_{m})(a)$. If $m$ is even, then, by (5.1), we have
@@
since $E(a)=b$ is an upper Christoffel word, and since $D,\widetilde{G}$ preserve upper Christoffel words, we obtain that $V_{m}$ is an upper Christoffel word. ∎
@@
Let $N=\sum_{1\leq i\leq m}d_{i}q_{i-1}$ be a greedy representation. Then $V_{m}(d_{1},\ldots,d_{m})$ is a Christoffel word if and only if the sequence $d_{1},\ldots,d_{m}$ is alternating. Precisely, for $m\geq 1$, the word $V_{m}(b_{1},0,b_{3},0,\ldots)$ is an upper Christoffel word, and the word $V_{m}(0,b_{2},0,b_{4},\ldots)$ is a lower Christoffel word.
Proof.
@@
Consider now the alternating sequence beginning by $b_{1}$, namely: $d_{1}=b_{1},d_{2}=0,\ldots$. Then $V_{1}=ba^{b_{1}},V_{2}=V_{1}^{b_{2}}a$ both begin by $b$, and therefore all $V_{i}$, $i\geq 1$, begin by $b$. ∎
Throughout the paper $\widetilde{x}$} denotes the reversal (mirror image) of the word $x$. A palindrome is a word equal to its reversal. The empty word is a palindrome. Recall that each proper lower Christoffel word $w$ has the factorization $w=apb$, where $p$ is a palindrome (called a central word), and that then the corresponding upper Christoffel word is $\widetilde{w}=bpa$, which is a conjugate of $w$. A standard word is $a$ or $b$, or a word of the from pabpab or pbapba for some central word $p$; it is known that standard words are obtained from $a,b$ by applying the endomorphisms in the submonoid generated by $G$ and $D$; moreover, each Christoffel class contains exactly two standard words pabpab and pbapba, where $p$ is the corresponding central palindrome. See [L, Subsection 2.2.1], [Re].
@@
Suppose that $m$ is even. Then by (5.1),
@@
Hence $M_{m}$ is a standard word. Since it ends by $b$, we have $M_{m}=pab$. ∎
We may derive a result, which is equivalent to a result previously obtained by Frid, [Fi] Corollary 1, and which is a noncommutative version of the Ostrowski representation.
@@
Let $N=0,\ldots,q_{m}-2$ be an integer with legal Ostrowski representation $N=\sum_{1\leq i\leq m}d_{i}q_{i-1}$. Then the prefix of length $N$ of the central palindrome $p$ is
@@
Proof.
The element $h=g^{-1}$, appearing in the proof of Theorem 7.3, is by (7.3) and (7.2) equal to
@@
In particular, $h$ is in $A^{*}$. Because of the inequalities (3.3), the word $V_{m}$ is in $A^{*}$, and $M_{m}$ is in $A^{*}$ too. Moreover, $h$ is of length $N$, by a calculation in the proof of Theorem 7.3; hence $|h|<q_{m}=|M_{m}|$. Thus, by (7.4) and by Lemma 4.1, $h$ is a prefix of $M_{m}$. Since $M_{m}=pab$ or $M_{m}=pba$ is of length $q_{m}$, we get that $h$ is a prefix of $p$. ∎
@@
This lemma could be deduced from a result of de Luca and Mignosi [dLM, Prop. 7]. We give an independent proof. See also Lemma 8.7 below.
Proof.
By stability, it is enough to prove this result for $i=m$. Suppose first that $b_{1}=c_{1}\geq 1$. We have by Lemma 7.1 and (5.1),
@@
By induction on the sum of the $a_{i}$, $u$ is a palindrome. Hence $G(u)a$ is a palindrome [Re, Lemma 4.1.4].
Suppose now that $b_{1}=0$, that is $a_{1}=1$. Then, since $V_{0}=a,V_{1}=b$, the sequence of words $V_{i},i=1,\ldots,m,$ is obtained from a shorter sequence, to which one applies $E$. We may therefore conclude by induction on $m$. ∎
@@
This result is not new: it is a consequence for example of the fact that the Burrows-Wheeler tableau of a Christoffel word (and even each perfectly clustering word) has a central symmetry (Theorem 4.3 of Simpson and Puglisi [SP]).
8. Borders of conjugates of Christoffel words

### v5

@@ Abstract
1 Introduction
In an article published in 1875 Christoffel (1875), Christoffel introduced a class of words on a binary alphabet, which now bear his name (following Berstel Berstel (1990)); they were shortly after rediscovered by Smith Smith (1876). Independently, in his 1880 work on minima of quadratic functions Markoff (1879, 1880), Markoff used these words to construct certain quadratic forms, which satisfy sharp inequalities relating their minima and their discriminant. Markoff was certainly unaware of the work of Christoffel. The relation between the work of Markoff and the Christoffel words was explicitly known to Frobenius Frobenius (1913), who also formulated the famous conjecture on the Markoff numbers (see Aigner (2013)). The theory of these words may be found in several books: Fogg (2002); Lothaire (2002); Berstel et al. (2009); Aigner (2013); Reutenauer (2019).
@@ 1 Introduction
1. They coincide with the elements of the free group with two generators subject to the following conditions: they are positive (that is, with no inverted letter), they are cyclically reduced, and they are part of a basis of this free group; see Osborne and Zieschang (1981); Kassel and Reutenauer (2007).
@@
2. They are the “perfectly clustering words” on a two-letter alphabet; this means that the last column of the Burrows-Wheeler tableau of such a word [^1] This tableau was defined by Burrows and Wheeler in the theory of data compression, see Restivo and Rosone (2011) Section 4., whose rows are the lexicographically sorted conjugates, is decreasing; see Mantaci et al. (2003); Ferenczi and Zamboni (2013).
@@
3. They constitute the finitary version of the Sturmian (infinite) words, which are obtained by discretizing straight lines in the plane, and which are characterized by the property that for each $n$, they have exactly $n+1$ factors of length $n$. For example, a word is a conjugate of a Christoffel word if and only if all its conjugates are factors of a Sturmian word; equivalently, this word (of length $n$ say) is primitive and has exactly n1n-1 circular factors of length n2n-2; see Lothaire (2002), (Reutenauer, 2019, Theorem 15.3.1).
@@
4. Besides the Christoffel words, which encode the Markoff forms and their minima, their conjugates correspond to the “small values” of these quadratic forms; see Reutenauer (2021).
@@
It is well known that Christoffel words are parametrized by nonnegative rational numbers. In the present article we first introduce a parametrization of the conjugates of Christoffel words, which is a finitary version of Bugeaud and Laurent (2023). This parametrization is based on the integer Ostrowski numeration system. It generalizes a construction which is widely used in the theory of Sturmian words, following Rauzy Rauzy (1985) (the “Rauzy rules”), and de Luca and Mignosi de Luca (1997) (the “standard words”). The construction is given in (11). Theorem 7.5 states that the whole conjugation class is constructed, and that it is independent of the chosen Ostrowski representation. As a corollary we obtain a result of Frid Frid (2018), which states (in some equivalent formulation) that the prefixes of a standard word are parametrized by legal Ostrowski representations; see Corollary 7.11, which appears as a noncommutative lifting of the Ostrowski numeration system.
@@
In Section 8, we study the borders (a prefix which is also a suffix) of conjugates of Christoffel words. It is well known that the length of the longest border of a word and its smallest period are simply related: their sum is the length of the word. The study of periods in words is an important matter in combinatorics on words, in particular in the theory of Sturmian words: it is known that each finite Sturmian word has a nontrivial proper period, except precisely the Christoffel words. We thus focus on conjugates of Christoffel words, and determine their longest borders. Our parametrization of the conjugates allows us to give precise statements on the form of these borders. In particular, they are themselves conjugates, or a power of them (Theorem 8.1, Corollaries 8.2 and 8.3). Note that smallest periods of conjugates of Christoffel words have been previously computed by Lapointe Lapointe (2017), and applied by her to the determination of normal forms, thereby allowing her to characterize conjugates within the class of Sturmian words. The set of smallest periods is also studied in Hegedüs and Nagy (2016) and Currie and Saari (2009).
@@
In Section 9, we give an application of our methods to notions and results due to Epifanio, Frougny, Gabriele, Mignosi and Shallit Epifanio et al. (2007, 2012). The result of Frid, once formulated for the so-called “lazy” Ostrowski representation (a notion introduced by these authors), may be translated into a result on the paths of a certain graph, called the “compact graph”; it states that each suffix of the central word corresponding to a Christoffel word is the label of a unique path, starting from the origin, in this graph. By specializing to lengths, one obtains the “Sturmian graph”: this graph has the property that each integer from 0 to the length of the central word is the label of a unique path; see Corollaries 9.3 and 9.5, due to Epifanio et al. (2007, 2012). As a consequence, we obtain the new result that these two graphs are naturally embedded in the tree of central words and in the Stern-Brocot tree; see Corollary 9.8, for the proof of which we use the “iterated palindromisation” of Aldo de Luca.
@@
Note that a link between lazy representations and periods of words was already established by Gabric, Rampersad and Shallit Gabric et al. (2021): they determine the set of periods of each prefix of length $n$ of a characteristic Sturmian (infinite) word and they show that the cardinality of this set is equal to the sum of the digits in the lazy representation of $n$.
@@ 2 Continuant polynomials
for any $k\geq 1$ (right recursion formula). It is customary to drop the index $k$ and to write $K(n_{1},\ldots,n_{k})$ for $K_{k}(n_{1},\ldots,n_{k})$, and in particular $K()=1$. One has (for example Cohn (1985) p. 116)
$P(n_{1})\cdots P(n_{k})=\left(\begin{array}[]{cc}K(n_{1},\ldots,n_{k})&K(n_{1},\ldots,n_{k-1})\\ K(n_{2},\ldots,n_{k})&K(n_{2},\ldots,n_{k-1})\end{array}\right),$ (1)
@@
It follows also, by transposing the product, and using the symmetry of the matrices $P(n)$, that we have $K(n_{1},\ldots,n_{k})=K(n_{k},\ldots,n_{1})$.
@@
$K(n_{1},\ldots,n_{k})=K(n_{1}-1,n_{2},\ldots,n_{k})+K(n_{2},\ldots,n_{k}),$ (2)
@@ 3 Ostrowski numeration
The representation (4) is called legal if one has the inequalities
$\forall i\geq 1,0\leq d_{i}\leq b_{i}.$ (5)
Among the legal representations, we distinguish two of them. We say that the representation (4) is greedy if it is legal and if the following condition is satisfied
$\forall i\geq 2,d_{i}=b_{i}\Rightarrow d_{i-1}=0.$ (6)
We say that the representation (4) is lazy if it is legal and if, with $k=\max\{i\mid d_{i}\neq 0\}$,
$\forall i,2\leq i\leq k,d_{i}=0\Rightarrow d_{i-1}=b_{i-1}.$ (7)
Proposition 3.1
@@
The existence of a representation (4) is implicit in Ostrowski’s article (Ostrowski, 1922, p.178); (i) is stated in Dupain (1979) p.83, and proved by Fraenkel, (Fraenkel, 1985, Theorem 3) (see also (Allouche and Shallit, 2003, Theorem 3.9.1) for a proof). Lazy Ostrowski representations were introduced by Epifanio, Frougny, Gabriele, Mignosi and Shallit in Epifanio et al. (2012); (ii) follows from their work.
@@
Lemma 3.2
Let $\sum_{j=1}^{m}d_{j}q_{j-1}$ be a greedy Ostrowski representation. Then, the inequality
$\sum_{j=1}^{m}(b_{j}-d_{j})q_{j-1}\leq q_{m-1}-1$
holds if and only if $d_{m}=b_{m}$ and the sequence $d_{i},i=1,\ldots,m$, is alternating.
Proof 3.3.
By Proposition 3.1 (ii), $\sum_{j=1}^{m}b_{j}q_{j-1}=q_{m}+q_{m-1}-2$ since the left-hand side is a lazy representation, and is necessarily the largest one. Thus the inequality of the lemma is equivalent to $\sum_{j=1}^{m}d_{j}q_{j-1}\geq q_{m}-1$. By the proposition again, part (i) this time, this inequality is equivalent to the fact that the left-hand side is the unique greedy representation of $q_{m}-1$. But by Lemma 10.1 (i), this unique representation is the alternating one, with $d_{m}=b_{m}$.
4 Conjugation
@@
Since $C^{n}$ is the identity on the words of length $n$, we have $C^{|g|}(u)=C^{r}(u)$, and this implies the last assertion.
5 Morphisms
@@
$\left(\begin{array}[]{cc}|f(g)|_{a}\\ |f(g)|_{b}\end{array}\right)=M(f)\left(\begin{array}[]{cc}|g|_{a}\\ |g|_{b}\end{array}\right).$ (9)
@@ 6 Christoffel words
The conjugation class of some Christoffel word is called a Christoffel class. It is known that in each Christoffel class, there is exactly one lower, and one upper, Christoffel word. See Lothaire (2002); Rauzy (1985) for this and other properties of these words.
@@
7 Construction of the conjugates of a Christoffel word
@@
Following Bugeaud and Laurent (2023), given a sequence of integers $d_{1},\ldots,d_{m}$ in $\mathbb{Z}$}, we define the following sequence $V_{i}=V_{i}(d_{1},\ldots,d_{m})$, of elements of $F(A)$, by
@@
The last one follows from the injectivity of the morphisms $\pi(i,j)$ (because they extend to automorphisms of the free group), hence of their product.
@@
Thus by (1) $|V_{m}|_{a}=K(b_{1},\ldots,b_{m})$ and $|V_{m}|_{b}=K(b_{2},\ldots,b_{m})$. We have $b_{i}=a_{i}$, except for $i=1$, where $b_{1}=a_{1}-1$. Thus (i) follows, using (2) for the third formula, and (ii) follows at once.
If the sequence $d_{1},\ldots,d_{m}$ satisfies the inequalities (5), then the exponents in (11) are all nonnegative, therefore $V_{i}(d_{1},\ldots,d_{m})\in A^{*}$. Define
$M_{m}=V_{m}(0,\ldots,0)=\pi(b_{1},0)\circ\cdots\circ\pi(b_{m},0)(a),$ (12)
@@
The $A^{*}$-conjugation class of $M_{m}$ is equal to the set of all $V_{m}(d_{1},\ldots,d_{m})$, for all sequences $d_{1},\ldots,d_{m}$ satisfying (5) and precisely $V_{m}=C^{N}(M_{m})$, with $N$ as above. This class contains the two Christoffel words of Slope $S=[0,a_{1},\ldots,a_{m}]$.
@@
$a^{-d_{1}}(\pi(b_{1},0)(a^{-d_{2}}))\cdots(\pi(b_{1},0)\circ\cdots\circ\pi(b_{m-1},0)(a^{-d_{m}})).$ (13)
@@
$V_{m}=h^{-1}M_{m}h$ (14)
and, by (10) and (9), the algebraic length of $h$ is equal to
@@
by (2), since $b_{i}=a_{i}$ if $i\geq 2$, and $b_{1}=a_{1}-1$.
If the sequence $d_{1},\ldots,d_{m}$ satisfies (5), then $V_{m}$ is in $A^{*}$, and by Lemma 4.1, $V_{m}=C^{N}(M_{m})$, thus $V_{m}$ is in the conjugation class of $M_{m}$. Conversely, each element of this class appears, since $M_{m}$ is of length $K(a_{1},\ldots,a_{m})$ (Lemma 7.3), and since, by Proposition 3.1, each $N=0,\ldots,K(a_{1},\ldots,a_{m})-1$ has an Ostrowski representation satisfying (5).
We show now that the class contains a Christoffel word. Consider the sequence $d_{1},\ldots,d_{m}$ defined by $d_{m}=b_{m},d_{m-1}=0,d_{m-2}=b_{m-2}$, and so on, depending on the parity of $m$. The corresponding element $V_{m}$ is in $A^{*}$, and is equal to $\cdots\pi(0,b_{m-2})\circ\pi(b_{m-1},0)\circ\pi(0,b_{m})(a)$. If $m$ is even, then, by (8), we have
@@
since $E(a)=b$ is an upper Christoffel word, and since $D,\widetilde{G}$ preserve upper Christoffel words, we obtain that $V_{m}$ is an upper Christoffel word.
@@
Let $N=\sum_{1\leq i\leq m}d_{i}q_{i-1}$ be a greedy representation. Then $V_{m}(d_{1},\ldots,d_{m})$ is a Christoffel word if and only if the sequence $d_{1},\ldots,d_{m}$ is alternating. Said more precisely, for $m\geq 1$, the word $V_{m}(b_{1},0,b_{3},0,\ldots)$ is an upper Christoffel word, and the word $V_{m}(0,b_{2},0,b_{4},\ldots)$ is a lower Christoffel word.
Proof 7.8.
@@
Consider now the alternating sequence beginning by $b_{1}$, namely: $d_{1}=b_{1},d_{2}=0,\ldots$. Then $V_{1}=ba^{b_{1}},V_{2}=V_{1}^{b_{2}}a$ both begin by $b$, and therefore all $V_{i}$, $i\geq 1$, begin by $b$.
Throughout the paper $\widetilde{x}$} denotes the reversal (mirror image) of the word $x$. A palindrome is a word equal to its reversal. The empty word is a palindrome. Recall that each proper lower Christoffel word $w$ has the factorization $w=apb$, where $p$ is a palindrome (called a central word), and that then the corresponding upper Christoffel word is $\widetilde{w}=bpa$, which is a conjugate of $w$. A standard word is $a$ or $b$, or a word of the from pabpab or pbapba for some central word $p$; it is known that standard words are obtained from $a,b$ by applying the endomorphisms in the submonoid generated by $G$ and $D$; moreover, each Christoffel class contains exactly two standard words pabpab and pbapba, where $p$ is the corresponding central palindrome. See (Lothaire, 2002, Subsection 2.2.1), Reutenauer (2019).
@@
Suppose that $m$ is even. Then by (8),
@@
Hence $M_{m}$ is a standard word. Since it ends by $b$, we have $M_{m}=pab$.
We may derive a result, which is equivalent to a result previously obtained by Frid, Frid (2018) Corollary 1, and which is a noncommutative version of the Ostrowski representation.
@@
Let $N=0,\ldots,q_{m}-2$ be an integer whose legal Ostrowski representation is given by $N=\sum_{1\leq i\leq m}d_{i}q_{i-1}$. Then the prefix of length $N$ of the central palindrome $p$ is
@@
Proof 7.12.
The element $h=g^{-1}$, appearing in the proof of Theorem 7.5, is by (13) and (12) equal to
@@
In particular, $h$ is in $A^{*}$. Because of the inequalities (5), the word $V_{m}$ is in $A^{*}$, and $M_{m}$ is in $A^{*}$ too. Moreover, $h$ is of length $N$, by a calculation in the proof of Theorem 7.5; hence $|h|<q_{m}=|M_{m}|$. Thus, by (14) and by Lemma 4.1, $h$ is a prefix of $M_{m}$. Since $M_{m}=pab$ or $M_{m}=pba$ is of length $q_{m}$, we get that $h$ is a prefix of $p$.
@@
This lemma could be deduced from a result of de Luca and Mignosi (de Luca and Mignosi, 1994, Prop. 7). We give an independent proof. See also Lemma 8.11 below.
Proof 7.14.
By stability, it is enough to prove this result for $i=m$. Suppose first that $b_{1}=c_{1}\geq 1$. We have by Lemma 7.1 and (8),
@@
By induction on the sum of the $a_{i}$, $u$ is a palindrome. Hence $G(u)a$ is a palindrome (Reutenauer, 2019, Lemma 4.1.4).
Suppose now that $b_{1}=0$, that is $a_{1}=1$. Then, since $V_{0}=a,V_{1}=b$, the sequence of words $V_{i},i=1,\ldots,m,$ is obtained from a shorter sequence, to which one applies $E$. We may therefore conclude by induction on $m$.
@@
This result is not new: it is a consequence for example of the fact that the Burrows-Wheeler tableau of a Christoffel word (and even each perfectly clustering word) has a central symmetry (Theorem 4.3 of Simpson and Puglisi Simpson and Puglisi (2008)).
8 Borders of conjugates of Christoffel words